# Beyond Redundancy: Toward Agile Resilience in Optical Networks to Overcome Unpredictable Disasters

### v1

@@ I Introduction
Network resilience is a fundamental requirement for ensuring service continuity for society and businesses. Even when infrastructure is damaged by natural disasters such as earthquakes, tsunamis, or floods, network services are expected to remain available. Conventional approaches to achieving resilience include (a) enhancing reliability through redundancy, (b) securing critical communications via congestion control and prioritized routing, and (c) enabling rapid restoration of services in accordance with predefined protocols [1]. These approaches typically rely on estimating disaster impact and service demand from demographic and facility locations; the importance of each facility is identified along with resource availability as key inputs for network design and disaster response. These approaches implicitly assume that traffic demand, facility location, and disaster scenarios can be predicted with reasonable accuracy. For instance, traffic demand, largely driven by human-to-human communication, has historically been inferred from demographic data, which are relatively stable and predictable. Network facility locations also tend to remain unchanged over decades of operation, while natural disasters have been predictable to some extent based on historical records. Under such conditions, these assumptions were regarded as valid.
@@
In recent years, however, the dominant type of service, climatic conditions, and geopolitical risks have undergone significant change, undermining the validity of the conventional assumptions of predictable traffic demand, facility locations, and disaster effects. Climate change due to global warming has led to unprecedented events, including floods and wildfires that cannot be explained by historical data [2]. In parallel, growing global awareness of potential risks facing infrastructure, including those related to military activities, which can co-occur in multiple locations [3], necessitates considering human-induced disruptions in addition to natural disasters. Such activites take into account redundancy-based disaster countermeasures and attempt to compromise infrastructure by exploiting the weakest points or targeting multiple locations simultaneously. Therefore, redundancy measures such as backup paths alone are insufficient to ensure resilience. Another significant change is that machine-to-machine communications have emerged as the dominant form of traffic [4]. Their demand is determined by the location of data centers hosting vast computational resources rather than by demographic data. While the prevailing trend once favored the construction of a few massive data centers, floor space and power supply constraints are now driving a shift toward a distributed data center architecture composed of multiple mid-sized data centers. As services increasingly depend on computing resources, resilience must be extended beyond networking to include computing infrastructures as well. At the same time, legal frameworks around the world are rapidly evolving to impose stricter data localization requirements and cross-border transfer restrictions, thereby generating regulatory uncertainty, which increase the difficulty of long-term infrastructure planning and investment. These emerging factors highlight the necessity of rethinking resilience approaches that rely on assumptions of predictability.
@@
To address the increasing unpredictability and scale of disasters, resilience must encompass not only static robustness within a single domain but also dynamic agility across multiple domains. Implementing comprehensive redundancy to cover all possible disasters is impractical from a cost perspective; therefore, redundancy is typically applied selectively based on disaster predictions. However, as disaster unpredictability continue to increase, the effectiveness of such pre-disaster countermeasures diminishes, and the need for rapid post-disaster recovery becomes more critical. In such cases, it is essential to aggregate available resources that remain unaffected by the disaster and dynamically interconnect them to restore services quickly. Moreover, as the scale of disasters grows, there is no guarantee that a single operator can secure the necessary resources for rapid recovery. In fact, Japanese telecommunication operators jointly launched a new cooperative framework to share facilities for rapid restoration in December 2024 [5]. We note that service continuity requires not only network resources but also computing resources. Consequently, leveraging resources distributed across multiple operators and layers, rather than relying solely on a single operator or a single layer, can significantly reduce the time required for service restoration. This paradigm shift requires us to develop agile resilience capabilities, enabling rapid adaptation to unforeseen failures and dynamic resource reallocation across multiple layers and operators.
@@
Although extensive studies have explored resilience techniques such as redundancy in optical networks and replication in distributed systems, these approaches remain limited because they either rely on pre-provisioned resources or operate within a single provider or layer, making them insufficient for achieving full agile resilience against large-scale or unpredictable disasters. Research on ring protection for dense wavelength division multiplexing (DWDM) transport [6] has been conducted for many years and provides high resilience with very short switching times, while multi-layer recovery has also been studied extensively [7]. However, both approaches depend on securing backup paths and equipment in advance, which works for predictable failures but is less effective for unforeseen situations. Similarly, considerable research on large-scale distributed service design has focused on commercial systems with embedded fault-tolerance mechanisms. The Google File System [8], for example, partitions files into chunks replicated across servers; in cloud environments, region replication establishes redundancy across geographically dispersed sites. Yet such methods usually assume users can access an alternative region after a large-scale disaster. Optical Spectrum as a Service (OSaaS) is a model that shares fiber among multiple operators [9]. It has been shown to optimize quality of transmission (QoT) using the channel-proving method; however, it is challenging to optimize line systems that degrade during disasters because the QoT optimization approach of OSaaS is based solely on bit error rate (BER) counts from transceivers.
@@ II-A Example Disaster Scenario
Operator A operates data centers in suburban area X and an optical network in an urban area (Fig. 1). They store confidential information (such as financial, medical, and health data) in databases, and use computing resources (e.g., CPUs and GPUs) to provide critical services primarily to urban users. Operator B also operates an optical network in the same urban area and runs data centers in suburban area Y, located some distance away from suburb X. Both operators own optical transmission equipment in the urban area and share several Points of Presence (PoPs).
@@
We consider an unexpected, large-scale natural disaster, such as a massive earthquake or tsunami, that causes widespread damage across Suburb X, resulting in power outages to data centers and optical transmission equipment for a period of 8 to 24 hours. Operator A’s pre-disaster preparedness measures prevent direct damage to the data centers and optical transmission equipment. However, Suburb X’s regional power grid or generation facilities suffer severe damage and stop working. To cope with such power outages, Operator A had deployed emergency backup power generators. The backup generators supply power to the facilities immediately after the power outage. However, following the Tier 1 and Tier 2 requirements of ANSI TIA-942, the stored fuel will run out in 8 to 24 hours.
As an emergency measure, Operator A temporarily leases optical transmission equipment and data center resources in Suburb Y, where the disaster impact is less severe, and tries to migrate critical services. While backup power is available, Operator A establishes a high-capacity optical wavelength path between Suburb X and Suburb Y by connecting its own optical transmission equipment to Operator B’s. Using the newly established wavelength paths, Operator A migrates the necessary computing resources, including databases, from Suburb X to Suburb Y, enabling continued service from data centers in Suburb Y. Because large-scale disasters simultaneously reduce available infrastructure capacity and increase communication demand for response and recovery operations, the rapidly provisioning of new high-capacity optical wavelength paths is preferred to relying solely on existing IP routing links. By rapidly establishing new optical wavelength paths over leased facilities and migrating computing resources, Operator A can enhance service continuity under catastrophic conditions, thereby improving its resilience.
@@ II-B Requirements and Challenges
To achieve agile resilience for rapid service migration, the followings are requirements:
  • (1) Rapid establishment of optical wavelength paths across borrowed infrastructure: Operators must be able to provision optical paths quickly using facilities beyond their own administrative domain.
@@
Accurate estimation of Quality of Transmission (QoT) is required to select appropriate modulation formats and launch powers for new wavelength paths. QoT depends on the characteristics of system components such as fibers, amplifiers, filters, and transceivers. Under normal conditions, these parameters (e.g., attenuation coefficient, chromatic dispersion, and nonlinear coefficient) are known from prior measurements. In post-disaster scenarios, however, the characteristics of borrowed equipment are often unknown, and the disaster itself may have altered them. Rapid path establishment therefore requires quickly extracting these parameters. Because measurement instruments and field engineers are scarce during the recovery phase, methods that minimize dedicated instruments and on-site intervention are highly desirable.
@@ III Key enablers for agile resilience
This section reviews state-of-the-art technologies related to agile resilience. First, we introduce fast and low-cost methods for characterization of borrowed transmission facilities, including fibers, amplifiers, and transceivers. Next, we outline peripheral technologies associated with transponder virtualization.
@@ III-A Facilities characterization method
This subsection introduces digital longitudinal monitoring (DLM) for fiber characterization, optical line system (OLS) calibration for amplifier characterization, and transceiver noise estimation for transceiver characterization.
@@ III-A3 Transceiver noise estimation
This method estimates transceiver noise using only a single variable optical attenuator (VOA) [13]. Transceivers also affect signal quality, and previous studies have shown that their impact can be modeled as additive Gaussian noise, similar to amplifier noise. The amount of transceiver noise is obtained from the back-to-back BER-OSNR curve, which represents the relationship between input OSNR and pre-FEC BER in a measurement setup without fiber propagation. Traditional BER-OSNR measurements require multiple optical devices, such as optical filters, amplifiers, spectrum analyzers, and ASE noise sources. In contrast, by exploiting the additive nature of Gaussian noise and the internally generated noise induced by changes in input power, rather than externally injected ASE noise, the BER-OSNR curve can be estimated by using only a VOA. This approach significantly reduces the measurement cost of transceiver characterization.
@@ III-B1 Resource isolation and access control
When delegating a subset of transponder line ports to other operators, it is necessary to isolate the associated resources and allow access only to that portion within the NOS. Two approaches can be considered: OS-level virtualization using virtual machines and process-level separation using containers. Computing nodes inside a data center, such as servers, can employ both approaches because they have abundant computing resources. However, since transponders are primarily designed for communication rather than computation and have limited computing resources, the lightweight container-based approach is preferred. For example, Telecom Infra Project (TIP) had been developing Goldstone, a container-based NOS [14]. We can achieve resource isolation and access control by implementing a container dedicated to controlling and managing specific line ports.
@@ III-B2 Open Control Interfaces
A virtualized transponder must be controllable and manageable by the delegated operator as intended. If its control interface is vendor-specific and closed, the delegation would impose additional cost and delay on the operator. Conversely, providing standardized open interfaces allows automation through APIs and seamless integration into existing ecosystems, thereby contributing to rapid service recovery. In fact, both TIP and OpenROADM MSA have been discussing open interfaces for transponder control, where YANG is adopted as the data model and NETCONF as the transport protocol [15]. Such open interfaces are expected to facilitate the rapid establishment of optical wavelength paths by the delegated operator.
@@ IV Validation through field evaluations
Figure 2: Field-trial optical network connecting Dublin and Galway for data center applications. (a) Geographical extension of the infrastructure. (b) Use-case functional representation. The topology emulates a network with two types of connections: the long-haul connection between two data centers located in different areas of the country, and a short reach one aiming to serve the metro area of an urban center. (c) Detailed hardware setup. The multi-vendor infrastructure implements a bidirectional linear optical network composed of two line systems and three nodes. Dark fibers and amplification sites are deployed in field alongside the production network. Color variation indicates vendor differences.
@@
This section aims to experimentally prove the feasibility of agile resilience in disaggregated optical networks through a field trial using an OLS to connect sites roughly 300 kilometers apart and key enablers (Sec. III). The presented experiment assumes a post-disaster scenario (Sec. II) in which the deployed equipment is borrowed from another operator. Without dispatching personnel to each local amplification site, the physical characteristics of the fiber and amplifier were quantified and incorporated into a physical model to virtually represent the transmission environment. Using this calibrated model, the network was configured to carry traffic through the deployment of multiple optical lightpaths. Subsequently, computing resources were migrated using the newly established communication path. In this experiment, we tested the transfer of a database.
@@
Fig.2 depicts the field-trial optical testbed deployed between the two cities of Dublin and Galway, designed to validate data center applications (Fig.2-a). The layout reflects two representative connectivity scenarios: a long-haul link interconnecting data centers situated in different regions of the country, and a metro-scale segment serving the needs of an urban area (Fig.2-b). At the hardware level, the setup follows a multi-vendor strategy and realizes a bidirectional linear optical topology, incorporating two line systems and three intermediate nodes (Fig.2-c). The evaluation was performed over dark fibers with amplification sites deployed in the field alongside the operational production network, thereby providing realistic conditions for assessing the proposed solution.
@@
Each line system is defined by a specific ROADM-to-ROADM connection. The long-haul link consists of 5 fiber spans including 4 in-line amplifiers (ILAs), with a total distance of 280 km. In this case, the two terminals of this link are equipped with OSAs and ASE noise source to perform the OLS calibration procedure and fill spectrum for troubleshooting and in-service transmission operation. The short reach link is emulated in the laboratory using a single fiber spool of 50 km. The portion of WDM spectrum implementing the data traffic is created by means of transceivers supporting 800G and 400G bit rates. The 800G wavelengths are obtained from NEC SpectralWave WX-T 800G and Fujitsu 1FINITY T950, while the 400G ones use NEC SpectralWave Phoenix WX-T. NEC NOS on SpectralWave has been used for transponder virtualization testing. This commercial NOS, based on TIP Goldstone NOS and an open control interface, deploys a dedicated container for the management of each line port. The rest of the full C-band spectrum is filled with 50-GHz dummy channels on a 100-GHz grid.

### v2

@@ I Introduction
Network resilience is a fundamental requirement for ensuring service continuity for society and businesses. Even when infrastructure is damaged by natural disasters such as earthquakes, tsunamis, or floods, network services are expected to remain available. Conventional approaches to achieving resilience include (a) enhancing reliability through redundancy, (b) securing critical communications via congestion control and prioritized routing, and (c) enabling rapid restoration of services in accordance with predefined protocols [1]. These approaches typically rely on estimating disaster impact and service demand from demographic and facility locations; the importance of each facility is identified along with resource availability as key inputs for network design and disaster response. These approaches implicitly assume that traffic demand, facility location, and disaster scenarios can reasonably predictable. For instance, traffic demand, largely driven by human-to-human communication, has historically been inferred from demographic data, which are relatively stable and predictable. Network facility locations also tend to remain unchanged over decades of operation, while natural disasters have been predictable to some extent based on historical records. Under such conditions, these assumptions were regarded as valid.
@@
In recent years, however, the dominant type of service, climatic conditions, and geopolitical risks have undergone significant change, undermining the validity of the conventional assumptions of predictable traffic demand, facility locations, and disaster effects. Climate change due to global warming has led to unprecedented events, including floods and wildfires that cannot be explained by historical data [2]. In parallel, growing global awareness of potential risks facing infrastructure, including those related to military activities, which can co-occur in multiple locations [3], necessitates considering human-induced disruptions in addition to natural disasters. Such activities take into account redundancy-based disaster countermeasures and attempt to compromise infrastructure by exploiting the weakest points or targeting multiple locations simultaneously. Therefore, redundancy measures such as backup paths alone are insufficient to ensure resilience. Another significant change is that machine-to-machine communications have emerged as the dominant form of traffic. Their demand is determined by the location of data centers hosting vast computational resources rather than by demographic data. While the prevailing trend once favored the construction of a few massive data centers, floor space and power supply constraints are now driving a shift toward a distributed data center architecture composed of multiple mid-sized data centers. As services increasingly depend on computing resources, resilience must be extended beyond networking to include computing infrastructures as well. At the same time, legal frameworks around the world are rapidly evolving to impose stricter data localization requirements and cross-border transfer restrictions, thereby generating regulatory uncertainty, which increase the difficulty of long-term infrastructure planning and investment. These emerging factors highlight the necessity of rethinking resilience approaches that rely on assumptions of predictability.
@@
To address the increasing unpredictability and scale of disasters, resilience must encompass not only static robustness within a single domain but also dynamic agility across multiple domains. Implementing comprehensive redundancy to cover all possible disasters is impractical from a cost perspective; therefore, redundancy is typically applied selectively based on disaster predictions. However, as disaster unpredictability continues to increase, the effectiveness of such pre-disaster countermeasures diminishes, and the need for rapid post-disaster recovery becomes more critical. In such cases, it is essential to aggregate available resources that remain unaffected by the disaster and dynamically interconnect them to restore services quickly. Moreover, as the scale of disasters grows, there is no guarantee that a single operator can secure the necessary resources for rapid recovery. In fact, Japanese telecommunication operators jointly launched the new cooperative framework to share facilities for rapid restoration in December 2024. This was further demonstrated in March 2025 through a field drill where four mobile operators installed base-station equipment on the cable-laying vessel of a Japanese operator to provide wireless coverage from the sea [4]. This exercise served as a concrete proof-of-concept of cross-carrier coordination and asset sharing for emergency network restoration. Similarly, in Australia, Telstra has collaborated with the government and other carriers to simulate temporary roaming on its network for users affected by outages [5]. We also note that service continuity requires not only network resources but also computing resources. Consequently, leveraging resources distributed across multiple operators and layers, rather than relying solely on a single operator or a single layer, can significantly reduce the time required for service restoration. This paradigm shift requires us to develop agile resilience capabilities, enabling rapid adaptation to unforeseen failures and dynamic resource reallocation across multiple layers and operators.
@@
Although extensive studies have explored resilience techniques such as redundancy in optical networks and replication in distributed systems, these approaches remain limited because they either rely on pre-provisioned resources or operate within a single provider or layer, making them insufficient for achieving full agile resilience against large-scale or unpredictable disasters. Research on ring protection for dense wavelength division multiplexing (DWDM) transport [6] has been conducted for many years and provides high resilience with very short switching times, while multi-layer recovery has also been studied extensively [7]. However, both approaches depend on securing backup paths and equipment in advance, which works for predictable failures but is less effective for unforeseen situations. Similarly, considerable research on large-scale distributed service design has focused on commercial systems with embedded fault-tolerance mechanisms. The Google File System [8], for example, partitions files into chunks replicated across servers; in cloud environments, region replication establishes redundancy across geographically dispersed sites. Yet such methods usually assume users can access an alternative region after a large-scale disaster.
@@ II-A Example Disaster Scenario
Operator A operates data centers in suburban area X and an optical network in an urban area (Fig. 1). They store confidential information (such as financial, medical, and health data) in databases, and use computing resources (e.g., CPUs and GPUs) to provide critical services primarily to urban users. Operator B also operates an optical network in the same urban area and provides critical services via data centers in suburban area Y. Both operators own optical transmission equipment in the urban area and share several Points of Presence (PoPs).
@@
We consider an unexpected, large-scale natural disaster, such as a massive earthquake or tsunami, that causes widespread damage across area X and area Y, resulting in power outages to data centers and optical transmission equipment within 72 to 96 hours. While pre-disaster preparedness measures prevent direct damage to the data centers and optical transmission equipment of both operators, the regional power grids or generation facilities suffer severe damage and cease operation. To mitigate such power outages, both operators have deployed emergency backup power generators. The backup generators supply power to the facilities immediately after the power outage. However, following the Tier 3 and Tier 4 requirements of ANSI TIA-942, the stored fuel will run out in 72 to 96 hours.
As an emergency measure, Operators A and B temporarily borrow data center Z resources in an area suffering relatively minor disaster impact, and cooperatively establish optical paths to the data center and attempt to migrate critical services from the affected areas while backup power is available. First, Operators A and B cooperatively establish a physical route between data center Z and the urban area by connecting their dark fibers. Next, they deploy spare optical transmission equipment (such as optical amplifiers and reconfigurable optical add & drop multiplexers (ROADMs)) — secured from their own central offices or maintenance depots, or rapidly provided from their procurement vendors — to the communication buildings along the fiber route to construct a temporary line system. Using this newly built line system, Operators A establishes optical paths between data center Z and area X. While backup power remains available, Operator A migrates essential computing resources, including databases, from area X to data center Z, ensuring service continuity from data center Z. Similarly, Operator A establishes additional optical paths between data center Z and area Y to migrate and maintain Operator B’s service as well. Because large-scale disasters simultaneously reduce available infrastructure capacity and increase communication demand for response and recovery operations, the rapid provisioning of high-capacity optical paths on newly built temporary infrastructure is preferred to relying solely on existing IP routing links. By rapidly establishing new optical paths over the ad-hoc infrastructure by leveraging spare assets and migrating computing resources, Operator A and B can enhance service continuity under catastrophic conditions, thereby improving their resilience.
@@ II-B Requirements and Challenges
To achieve agile resilience for rapid service migration, the following are requirements:
  • (1) Rapid establishment of optical paths across cooperatively built ad-hoc infrastructure: Operators must be able to quickly provision optical paths on a newly built line system that leverages spare assets, such as dark fibers and spare equipment, across multiple administrative domains.
@@
Accurate estimation of Quality of Transmission (QoT) is required to select appropriate modulation formats and launch powers for new wavelength paths. QoT depends on the characteristics of system components such as fibers, amplifiers, filters, and transceivers. Under normal conditions, these parameters (e.g., attenuation coefficient, chromatic dispersion, and nonlinear coefficient) are known from prior measurements. In the agile resilience scenarios considered here, a new temporary line system will be constructed using dark fibers and spare maintenance equipment; the characteristics of such unused resources are not monitored continuously, as they are typically measured only prior to use. Even if previous measurement records exist, physical characteristics may have changed due to aging or the impact of the disaster itself. Rapid path establishment therefore requires quick and accurate extraction of these parameters. Because measurement instruments and field personnel may be limited during the recovery phase, methods that minimize dedicated instruments and on-site intervention are highly desirable.
To address this challenge, we introduce Digital Longitudinal Monitoring (DLM) and Optical Line System (OLS) calibration. These technologies enable the rapid extraction of fiber and amplifier parameters while using only the transponders and built-in power meters at the path endpoints, eliminating the need for dispatching field personnel or employing dedicated measurement instruments.
@@ III Key enablers for agile resilience
This section reviews state-of-the-art technologies related to agile resilience. First, we introduce fast and low-cost methods for the characterization of borrowed transmission facilities, including fibers, amplifiers, and transceivers. Next, we outline peripheral technologies associated with transponder virtualization.
@@ III-A Facilities characterization method
This subsection introduces DLM for fiber characterization, OLS calibration for amplifier characterization, and transceiver noise estimation for transceiver characterization.
@@ III-A3 Transceiver noise estimation
This method estimates transceiver noise using only a single variable optical attenuator (VOA) [12]. Transceivers also affect signal quality, and previous studies have shown that their impact can be modeled as additive Gaussian noise, similar to amplifier noise [13]. The amount of transceiver noise is obtained from the back-to-back BER-OSNR curve, which represents the relationship between input OSNR and pre-FEC BER in a measurement setup without fiber propagation. Traditional BER-OSNR measurements require multiple optical devices, such as optical filters, amplifiers, spectrum analyzers, and ASE noise sources. In contrast, by exploiting the additive nature of Gaussian noise and the internally generated noise induced by changes in input power, rather than externally injected ASE noise, the BER-OSNR curve can be estimated by using only a VOA. This approach significantly reduces the measurement cost of transceiver characterization.
@@ III-B1 Resource isolation and access control
When delegating a subset of transponder line ports to other operators, it is necessary to isolate the associated resources and allow access only to that portion within the NOS. Two approaches can be considered: OS-level virtualization using virtual machines and process-level separation using containers. Computing nodes inside a data center, such as servers, can employ both approaches because they have abundant computing resources. However, since transponders are primarily designed for communication rather than computation and have limited computing resources, the lightweight container-based approach is advantageous. For example, Telecom Infra Project (TIP) had been developing Goldstone, a container-based NOS [14]. We can achieve resource isolation and access control by implementing a container dedicated to controlling and managing specific line ports.
@@ III-B2 Open Control Interfaces
A virtualized transponder must be controllable and manageable by the delegated operator as intended. If its control interface is vendor-specific and closed, the delegation would impose additional cost and delay on the operator. Conversely, providing standardized open interfaces allows automation through APIs and seamless integration into existing ecosystems, thereby contributing to rapid service recovery. In fact, both TIP and Open ROADM MSA have been discussing open interfaces for transponder control, where YANG is adopted as the data model and NETCONF as the transport protocol [15]. Such open interfaces are expected to facilitate the rapid establishment of optical paths by the delegated operator.
@@ IV Validation through field evaluations
Figure 2: Field-trial optical network connecting Dublin and Galway. (a) Geographical layout. (b) Use-case functional representation. The topology emulates a long-haul connection between two data centers and a short reach metro link. (c) Hardware setup. The multi-vendor, bidirectional network consisting of two line systems and three nodes. Dark fibers and amplification sites are deployed in the field alongside the production network. Color indicates vendor differences.
@@
This section aims to experimentally prove the feasibility of agile resilience in disaggregated optical networks through a field trial using an OLS to connect sites roughly 300 km apart and the key enablers (Sec. III). The presented experiment assumes a post-disaster scenario (Sec. II) in which the deployed equipment is borrowed from another operator. Without dispatching personnel to each local amplification site, the physical characteristics of the fiber and amplifier were quantified and incorporated into a physical model. Using this calibrated model, multiple optical paths were deployed to carry traffic. Subsequently, computing resources, specifically a database, were migrated using the newly established communication path.
IV-A Setup
@@
Fig.2 depicts the field-trial optical testbed deployed between the two cities of Dublin and Galway, designed to validate data center applications (Fig.2-a). The layout reflects two representative connectivity scenarios: a long-haul link interconnecting data centers situated in different regions of the country, and a metro-scale segment serving the needs of an urban area (Fig.2-b). The setup follows a multi-vendor strategy and realizes a bidirectional linear optical topology, incorporating two line systems and three intermediate nodes (Fig.2-c). The evaluation was performed over dark fibers with amplification sites deployed in the field alongside the operational production network, thereby providing realistic conditions for assessing the proposed solution.
@@
Each line system is defined by a specific ROADM-to-ROADM connection. The 280 km long-haul link consists of five fiber spans including four in-line amplifiers (ILAs). The two terminals of the long-haul link are equipped with OSAs and an ASE noise source to perform the OLS calibration procedure and fill spectrum for troubleshooting and in-service transmission operation. The short-reach link is emulated in the laboratory using a single fiber spool of 50 km. The portion of WDM spectrum implementing the data traffic is created by means of transceivers supporting 800G and 400G bit rates. The 800G wavelengths are obtained from NEC SpectralWave WX-T 800G and Fujitsu 1FINITY T950, while the 400G ones use NEC SpectralWave Phoenix WX-T. NEC NOS on SpectralWave has been used for transponder virtualization testing. This commercial NOS, based on TIP Goldstone NOS and an open control interface, deploys a dedicated container for the management of each line port. The rest of the full C-band spectrum is filled with 50-GHz dummy channels on a 100-GHz grid.
IV-B Procedure and time record